OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. 3x^2 - 4x - 8 = 0 A. {-1.7, 2.8} B. {-2.8, 2} C. {-1.1, 2.4} D. no real solution

OpenStudy (anonymous):

you have a equation in the form ax^2 + bx + c = 0. The quadratic equation is \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]. Just plug in a, b, and c in and simplify

OpenStudy (anonymous):

^_- huh

OpenStudy (anonymous):

In the equation 3x^2 - 4x - 8 a=3 b=-4 c=-8 just plug those numbers into the formula and simplify

OpenStudy (anonymous):

−4±-4^2-4 (-4)(3)(8) − − − − − − − √ 2a

OpenStudy (amistre64):

dont forget you a spot in the denominator

OpenStudy (amistre64):

the latex format doesnt copy paste to well does it :)

OpenStudy (anonymous):

no

OpenStudy (amistre64):

-b +- sqrt(b^2 - 4ac) ------------------- 2a that +- is notation for 2 separate formats; the are identical except for one has a + there and the other a - a +- b means: a+b and a-b

OpenStudy (anonymous):

And its -b in the front so it would be - (-4) in the beginning

OpenStudy (anonymous):

Ok you have both lost me

OpenStudy (anonymous):

\[\frac{-(-4) \pm \sqrt{(-4)^2 - 4(3)(-8)}}{2(3)}\]

OpenStudy (anonymous):

I just plugged a, b and c into the formula. All you have to do is simplify